# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v4

@@ 6.2 WiLDS Datasets
The WILDS Camelyon17 dataset consists of images of cancerous and non-cancerous cell tissue. We use the “train-unlabeled" split for pre-training ExPLoRA, and either use LoRA fine-tuning or linear probing on the training set of the labeled split. We report accuracy on the binary classification problem and compare with entries on the WILDS leaderboard which use unlabeled data. Our results in table 7 demonstrate improved performance over domain-specific methods as well as DinoV2, once again successfully bridging the domain gap.
@@
iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set, finding that this outperforms pre-training on the extra-unlabeled set. In table 8, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible overfitting due to a small domain gap. This is likely because natural image datasets such as ImageNet (Deng et al., 2009) used for pre-training DinoV2 already contain many images of animals.
@@ 7 Further Analysis
The key design choice of ExPLoRA is to constrain the trainable parameter space during pre-training via $\Theta({\mathcal{U}},r)$ in eq. 5, where ${\mathcal{U}}\subset{\mathcal{L}}$ layers of the ViT undergo full-rank training while the remaining frozen layers ${\mathcal{L}}\setminus{\mathcal{U}}$ receive low-rank updates. For parameter-efficiency, we wish to keep $|{\mathcal{U}}|\ll|{\mathcal{L}}|$ and make an informed choice of which layers to unfreeze based on their potential to improve representation learning during extended pre-training.
We conduct an investigation on 5 models using a sample of ${\mathcal{X}}_{D_{T}}$. These models are DinoV2, D-ExPLoRA-[L]-$r64$, SatMAE, MAE, and M-ExPLoRA-[L]-$r64$. We do the following analyses: (i) PCA to measure the mean and variance of eigenvalues of patch feature vectors for each ViT block, in figs. 3 and 4 (ii) linear probing for local or global information (Darcet et al., 2024) by training logistic regression classifiers on each block’s patch feature vectors, to predict either patch position (fig. 5) or image class (fig. 6).
Findings and Unfreezing Strategy for DinoV2: Our analysis reveals that the spectral properties of a block’s feature map (fig. 3) and the ability to retrieve local information from its output patch tokens (fig. 5) are correlated. The classification accuracy for position and the mean and variance of the principal eigenvalues peak in the middle-layers of the model, suggesting that the middle blocks capture fine-grained local properties of patches (e.g., texture, relative position). Meanwhile, deeper blocks focus on global semantic understanding, as shown by increased classification accuracy for image class prediction in fig. 6 and lower variance in feature map eigenvalues in fig. 4. Combined, these results suggest that unfreezing deeper layers, such as ${\mathcal{U}}=\{\texttt{L}\}$, allows the model to better capture global features without overfitting to local details of images of $T$. This is empirically confirmed in table 3, where linear probing accuracy correlates inversely with the mean eigenvalue of each block (i.e., block 23 > block 22 > block 0 > block 9). The attention maps in fig. 9 further support this, showing that deeper layers focus more clearly on central objects, while earlier layers (e.g., blocks 9, 10) exhibit diffuse attention patterns spread around the border.
@@
Thus, our analysis provides guidelines for practitioners to select which blocks to unfreeze based on the eigenvalue properties and classification accuracy patterns of different ViT layers, offering a systematic approach to constraining $\Theta({\mathcal{U}},r)$ when pre-training with ExPLoRA (eq. 5).
@@ 8 Conclusion and Discussion
For example, automated analysis of satellite imagery can inform social, economic, and environmental policies, but manual curation is expensive, and pre-training models on such data has significant costs, both environmental and otherwise (appendix D). ExPLoRA offers a more efficient way to create new foundation models for different visual domains via distilling knowledge from existing foundation models trained on natural images. This can sharply reduce costs, aid researchers and policymakers, and enable flexible downstream applications.
@@ Acknowledgements
This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via 2021-2011000004, NSF(#1651565), ARO (W911NF-21-1-0125), ONR (N00014-23-1-2159), CZ Biohub, HAI. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes not-withstanding any copyright annotation therein.
@@
We include supplementary material in the following sections. We provide additional contextualization and comparison with recent related work in appendix A, additional experimental results and ablations in appendix B, training and hyperparameter details in appendix C, and an overview of the environmental impact of our work in appendix D.
@@ A.1 Comparison with Geospatial Domain Adaptation
Recent work has explored both continual pre-training (GFM) (Mendieta et al., 2023) and parameter-efficient domain adaptation (GDA) (Scheibenreif et al., 2024) for satellite imagery. We compare these approaches with ExPLoRA in table 10.
@@
We also demonstrate ExPLoRA’s broader applicability through experiments on larger datasets (fMoW-RGB, fMoW-Sentinel, which have 400k-800k images vs 90k images in FireRisk (Shen et al., 2023), the largest dataset used in GDA) and domains beyond remote sensing (i.e. WiLDS). Our analysis in section 7 provides insights into block-wise information encoding, offering practitioners a systematic approach for block selection during extended pre-training– a unique feature not present in prior work.
@@ A.2 Comparison with Unsupervised Domain Adaptation
ExPLoRA’s setting in section 4 is different: we only require weights $W_{S}$ from unsupervised pre-training on $p_{S}({\mathbf{x}})$, without source data access or label set restrictions. This enables adaptation across wider domain shifts (e.g., ImageNet to multi-spectral satellite imagery, section 6.1.3). Thus, rather than competing with UDA methods, ExPLoRA can complement them by providing a better initialization than standard natural-image pre-training (as seen empirically in section B.7).
@@ Appendix B Additional Experimental Results
Below, we include further experimental results as a continuation of section 6.
@@ B.1 Results on Additional Downstream Datasets
We consider a land-cover classification dataset used in (Ayush et al., 2021), where each of 244,471 training and 55,529 validation images are paired with one of 66 land cover classes obtained by the USDA’s National Agricultural Imagery Program. In table 12, we first demonstrate similar performance between both natural-image backbones (rows 4 and 5), which surprisingly outperform SatMAE, which is pre-trained on fMoW-RGB. We use ExPLoRA to pre-train from DinoV2 to the training set of this dataset (without labels). Our results (row 6) demonstrate comparable performance, suggesting that for this dataset, domain-specific knowledge may not be highly relevant to successfully solve the task.
@@
The dataset contains 27,000 13-band satellite images of 10 classes (Helber et al., 2019), sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on this dataset’s training set, and instead use LoRA fine-tuning starting with the pre-trained weights learned in the last row of table 4. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table 13). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
@@ B.2 The Importance of Extended Pre-training
We address these questions in fig. 7, focusing on DinoV2 models running on NVIDIA-A4000 GPUs. We evaluate D-ExPLoRA-$\tt{[L]}$-$r64$ for different lengths of pre-training (50k, 100k, 150k, and 200k iterations), corresponding to 24, 48, 72, and 96 GPU-hours of extended pre-training respectively. Each checkpoint undergoes LoRA-$r8$ fine-tuning. We compare against three baselines: (i) Direct LoRA-$r8$ fine-tuning on DinoV2 weights (ii) Fine-tuning DinoV2 with block 24 unfrozen and LoRA-$r64$ (matching ExPLoRA’s parameter budget) (iii) Fine-tuning DinoV2 with blocks 0, 1, 23, 24 unfrozen and LoRA-$r64$ (55.8M parameters vs ExPLoRA’s 18.7M).
Results in fig. 7 demonstrate that ExPLoRA’s extended pre-training achieves a $\uparrow 1.0\%$ improvement in maximum fine-tuning accuracy within the same total compute budget (320 GPU hours). Notably, even increasing the parameter budget during fine-tuning fails to match this performance. While additional pre-training iterations beyond 100k improve initial fine-tuning accuracy, they have minimal impact on the final accuracy ceiling, highlighting ExPLoRA’s computational efficiency. Importantly, the total amount of compute invested in pre-training and fine-tuning ExPLoRA is still dwarfed by the compute spent for full pre-training from scratch, e.g., 960 GPU hours for SatMAE (table 1).
Lastly, we re-iterate that the benefits of extended pre-training become more pronounced as the domain gap increases. While fig. 7 demonstrates a 1% improvement on RGB satellite images, table 4 highlights a 8% improvement when using ExPLoRA + PEFT (row 8) over full fine-tuning MAE (row 1) on multi-spectral satellite images– a domain with substantial shift from natural images. ExPLoRA still requires less total compute (610 vs 770 GPU hours), demonstrating that its efficiency advantages scale with domain difficulty.
@@ B.3 Convergence and Data Efficiency
In fig. 8, we plot the linear-probing accuracy against the number of extended pre-training iterations for ExPLoRA (in blue). ExPLoRA improves quickly, requiring between 100-150k extended pre-training iterations to reach optimal performance. As discussed in section 6.1.2, unfreezing additional transformer blocks (in red) fails to achieve the same level of performance while requiring more parameters.
@@ B.4 Impact of MAE Decoder Rank
As outlined in section 5, we initialize the MAE decoder $g_{\psi}$ (with its ${\mathcal{L}}_{D}$ transformer layers) with pre-trained weights $W_{S}$ from He et al. (2022), keeping all decoder weights (except layer norm) frozen during extended pre-training on ${\mathcal{X}}_{T}$. We apply LoRA with rank $r^{\prime}$ to the $Q,V$ weights of the attention layers in the decoder ${\mathcal{L}}_{D}$, while unfreezing 1-2 blocks ${\mathcal{U}}$ in the ViT encoder ${\mathcal{L}}$ and applying LoRA with rank $r$ to the remaining layers ${\mathcal{L}}\setminus{\mathcal{U}}$ (algorithm 1).
We evaluate ExPLoRA with M-$\tt{[1,L]}$-$r64$ on fMoW-Sentinel, using a fixed encoder LoRA rank $r=64$, unfreezing blocks ${\mathcal{U}}=\{\tt{1,L}\}$, and varying the decoder rank $r^{\prime}$. We then fine-tune the resulting model with LoRA $r=8$ and measure the highest top 1 accuracy on the validation set of fMoW-Sentinel. Table 14 shows that increasing $r^{\prime}$ up to 32 improves fine-tuning performance, which then declines by $\downarrow 0.94\%$ for $r^{\prime}=64$. This suggests that balancing the unfrozen parameters between the ViT encoder $f_{\theta}$ (used for fine-tuning) and the MAE decoder $g_{\psi}$ (discarded post pre-training) is crucial. Larger $r^{\prime}$ may improve the decoder’s ability without benefiting the learned representations of $f$. This issue doesn’t arise in DinoV2, as the Dino-iBOT shared head is fully trained since it isn’t provided by Oquab et al. (2024).
@@ B.5 Impact of ViT backbone size
We also test the impact of the ViT backbone for ExPLoRA, varying the architecture for DinoV2 from ViT-B (86M, $\tt{L}=12$ layers, embedding dimension 768), ViT-L (303M parameters, $\tt{L}=24$ layers, embedding dimension 1024), and ViT-G (1100M parameters, $\tt{L}=40$ layers, embedding dimension 1280) for extended pre-training on fMoW-RGB. The ExPLoRA models we compare against are D-[12]-$r64$ for ViT-B, D-[24]-$r64$ for ViT-L, and D-[32]-$r32$ for ViT-G. We unfreeze the 12th, 24th, and 32nd layers for each of ViT-B, ViT-L, and ViT-G, picking these layers by extending the analysis from section 7 to ViT-B and ViT-G. We find that the 12th (last layer) for ViT-B and the 32nd (out of 40) layer for ViT-G output representations with low mean eigenvalues compared to other layers, thus presenting good candidates for unfreezing.
In table 15, we see that as expected, ViT-G performs the best, but is only $\uparrow 0.31\%$ better in top 1 accuracy compared to ViT-L, while using many more parameters. On the other hand, we see the highest impact for ExPLoRA on ViT-B, where the top 1 accuracy improves by $\uparrow 9.21\%$ over the original DinoV2 ViT-B. These results further demonstrate the effectiveness and efficiency of ExPLoRA as a powerful technique to create unsupervised foundation models for new visual domains.

### v5

@@ 6.2 WiLDS Datasets
The WILDS Camelyon17 dataset consists of images of cancerous and non-cancerous cell tissue. We use the “train-unlabeled" split for pre-training ExPLoRA, and either use LoRA fine-tuning or linear probing on the training set of the labeled split. We report accuracy on the binary classification problem and compare with entries on the WILDS leaderboard which use unlabeled data. Our results in table˜7 demonstrate improved performance over domain-specific methods as well as DinoV2, once again successfully bridging the domain gap.
@@
iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set, finding that this outperforms pre-training on the extra-unlabeled set. In table˜8, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible overfitting due to a small domain gap. This is likely because natural image datasets such as ImageNet (Deng et al., 2009) used for pre-training DinoV2 already contain many images of animals.
@@ 7 Further Analysis
The key design choice of ExPLoRA is to constrain the trainable parameter space during pre-training via $\Theta({\mathcal{U}},r)$ in eq.˜5, where $\mathcal{U}$$\subset{\mathcal{L}}$ layers of the ViT undergo full-rank training while the remaining frozen layers ${\mathcal{L}}\setminus{\mathcal{U}}$ receive low-rank updates. For parameter-efficiency, we wish to keep $|{\mathcal{U}}|\ll|{\mathcal{L}}|$ and make an informed choice of which layers to unfreeze based on their potential to improve representation learning during extended pre-training.
We conduct an investigation on 5 models using a sample of ${\mathcal{X}}_{D_{T}}$. These models are DinoV2, D-ExPLoRA-[L]-$r64$, SatMAE, MAE, and M-ExPLoRA-[L]-$r64$. We do the following analyses: (i) PCA to measure the mean and variance of eigenvalues of patch feature vectors for each ViT block, in figs.˜3 and 4 (ii) linear probing for local or global information (Darcet et al., 2024) by training logistic regression classifiers on each block’s patch feature vectors, to predict either patch position (fig.˜5) or image class (fig.˜6).
Findings and Unfreezing Strategy for DinoV2: Our analysis reveals that the spectral properties of a block’s feature map (fig.˜3) and the ability to retrieve local information from its output patch tokens (fig.˜5) are correlated. The classification accuracy for position and the mean and variance of the principal eigenvalues peak in the middle-layers of the model, suggesting that the middle blocks capture fine-grained local properties of patches (e.g., texture, relative position). Meanwhile, deeper blocks focus on global semantic understanding, as shown by increased classification accuracy for image class prediction in fig.˜6 and lower variance in feature map eigenvalues in fig.˜4. Combined, these results suggest that unfreezing deeper layers, such as ${\mathcal{U}}=\{\texttt{L}\}$, allows the model to better capture global features without overfitting to local details of images of $T$. This is empirically confirmed in table˜3, where linear probing accuracy correlates inversely with the mean eigenvalue of each block (i.e., block 23 > block 22 > block 0 > block 9). The attention maps in fig.˜9 further support this, showing that deeper layers focus more clearly on central objects, while earlier layers (e.g., blocks 9, 10) exhibit diffuse attention patterns spread around the border.
@@
Thus, our analysis provides guidelines for practitioners to select which blocks to unfreeze based on the eigenvalue properties and classification accuracy patterns of different ViT layers, offering a systematic approach to constraining $\Theta({\mathcal{U}},r)$ when pre-training with ExPLoRA (eq.˜5).
@@ 8 Conclusion and Discussion
For example, automated analysis of satellite imagery can inform social, economic, and environmental policies, but manual curation is expensive, and pre-training models on such data has significant costs, both environmental and otherwise (appendix˜D). ExPLoRA offers a more efficient way to create new foundation models for different visual domains via distilling knowledge from existing foundation models trained on natural images. This can sharply reduce costs, aid researchers and policymakers, and enable flexible downstream applications.
@@ Acknowledgements
This research is based upon work supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via 2021-2011000004, NSF(#1651565), ARO (W911NF-21-1-0125), ONR (N00014-23-1-2159), CZ Biohub, HAI. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@
We include supplementary material in the following sections. We provide additional contextualization and comparison with recent related work in appendix˜A, additional experimental results and ablations in appendix˜B, training and hyperparameter details in appendix˜C, and an overview of the environmental impact of our work in appendix˜D.
@@ A.1 Comparison with Geospatial Domain Adaptation
Recent work has explored both continual pre-training (GFM) (Mendieta et al., 2023) and parameter-efficient domain adaptation (GDA) (Scheibenreif et al., 2024) for satellite imagery. We compare these approaches with ExPLoRA in table˜10.
@@
We also demonstrate ExPLoRA’s broader applicability through experiments on larger datasets (fMoW-RGB, fMoW-Sentinel, which have 400k-800k images vs 90k images in FireRisk (Shen et al., 2023), the largest dataset used in GDA) and domains beyond remote sensing (i.e. WiLDS). Our analysis in section˜7 provides insights into block-wise information encoding, offering practitioners a systematic approach for block selection during extended pre-training– a unique feature not present in prior work.
@@ A.2 Comparison with Unsupervised Domain Adaptation
ExPLoRA’s setting in section˜4 is different: we only require weights $W_{S}$ from unsupervised pre-training on $p_{S}({\mathbf{x}})$, without source data access or label set restrictions. This enables adaptation across wider domain shifts (e.g., ImageNet to multi-spectral satellite imagery, section˜6.1.3). Thus, rather than competing with UDA methods, ExPLoRA can complement them by providing a better initialization than standard natural-image pre-training (as seen empirically in section˜B.7).
@@ Appendix B Additional Experimental Results
Below, we include further experimental results as a continuation of section˜6.
@@ B.1 Results on Additional Downstream Datasets
We consider a land-cover classification dataset used in (Ayush et al., 2021), where each of 244,471 training and 55,529 validation images are paired with one of 66 land cover classes obtained by the USDA’s National Agricultural Imagery Program. In table˜12, we first demonstrate similar performance between both natural-image backbones (rows 4 and 5), which surprisingly outperform SatMAE, which is pre-trained on fMoW-RGB. We use ExPLoRA to pre-train from DinoV2 to the training set of this dataset (without labels). Our results (row 6) demonstrate comparable performance, suggesting that for this dataset, domain-specific knowledge may not be highly relevant to successfully solve the task.
@@
The dataset contains 27,000 13-band satellite images of 10 classes (Helber et al., 2019), sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on this dataset’s training set, and instead use LoRA fine-tuning starting with the pre-trained weights learned in the last row of table˜4. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table˜13). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
@@ B.2 The Importance of Extended Pre-training
We address these questions in fig.˜7, focusing on DinoV2 models running on NVIDIA-A4000 GPUs. We evaluate D-ExPLoRA-$\tt{[L]}$-$r64$ for different lengths of pre-training (50k, 100k, 150k, and 200k iterations), corresponding to 24, 48, 72, and 96 GPU-hours of extended pre-training respectively. Each checkpoint undergoes LoRA-$r8$ fine-tuning. We compare against three baselines: (i) Direct LoRA-$r8$ fine-tuning on DinoV2 weights (ii) Fine-tuning DinoV2 with block 24 unfrozen and LoRA-$r64$ (matching ExPLoRA’s parameter budget) (iii) Fine-tuning DinoV2 with blocks 0, 1, 23, 24 unfrozen and LoRA-$r64$ (55.8M parameters vs ExPLoRA’s 18.7M).
Results in fig.˜7 demonstrate that ExPLoRA’s extended pre-training achieves a $\uparrow 1.0\%$ improvement in maximum fine-tuning accuracy within the same total compute budget (320 GPU hours). Notably, even increasing the parameter budget during fine-tuning fails to match this performance. While additional pre-training iterations beyond 100k improve initial fine-tuning accuracy, they have minimal impact on the final accuracy ceiling, highlighting ExPLoRA’s computational efficiency. Importantly, the total amount of compute invested in pre-training and fine-tuning ExPLoRA is still dwarfed by the compute spent for full pre-training from scratch, e.g., 960 GPU hours for SatMAE (table˜1).
Lastly, we re-iterate that the benefits of extended pre-training become more pronounced as the domain gap increases. While fig.˜7 demonstrates a 1% improvement on RGB satellite images, table˜4 highlights a 8% improvement when using ExPLoRA + PEFT (row 8) over full fine-tuning MAE (row 1) on multi-spectral satellite images– a domain with substantial shift from natural images. ExPLoRA still requires less total compute (610 vs 770 GPU hours), demonstrating that its efficiency advantages scale with domain difficulty.
@@ B.3 Convergence and Data Efficiency
In fig.˜8, we plot the linear-probing accuracy against the number of extended pre-training iterations for ExPLoRA (in blue). ExPLoRA improves quickly, requiring between 100-150k extended pre-training iterations to reach optimal performance. As discussed in section˜6.1.2, unfreezing additional transformer blocks (in red) fails to achieve the same level of performance while requiring more parameters.
@@ B.4 Impact of MAE Decoder Rank
As outlined in section˜5, we initialize the MAE decoder $g_{\psi}$ (with its ${\mathcal{L}}_{D}$ transformer layers) with pre-trained weights $W_{S}$ from He et al. (2022), keeping all decoder weights (except layer norm) frozen during extended pre-training on ${\mathcal{X}}_{T}$. We apply LoRA with rank $r^{\prime}$ to the $Q,V$ weights of the attention layers in the decoder ${\mathcal{L}}_{D}$, while unfreezing 1-2 blocks $\mathcal{U}$} in the ViT encoder $\mathcal{L}$} and applying LoRA with rank $r$ to the remaining layers ${\mathcal{L}}\setminus{\mathcal{U}}$ (algorithm˜1).
We evaluate ExPLoRA with M-$\tt{[1,L]}$-$r64$ on fMoW-Sentinel, using a fixed encoder LoRA rank $r=64$, unfreezing blocks ${\mathcal{U}}=\{\tt{1,L}\}$, and varying the decoder rank $r^{\prime}$. We then fine-tune the resulting model with LoRA $r=8$ and measure the highest top 1 accuracy on the validation set of fMoW-Sentinel. Table˜14 shows that increasing $r^{\prime}$ up to 32 improves fine-tuning performance, which then declines by $\downarrow 0.94\%$ for $r^{\prime}=64$. This suggests that balancing the unfrozen parameters between the ViT encoder $f_{\theta}$ (used for fine-tuning) and the MAE decoder $g_{\psi}$ (discarded post pre-training) is crucial. Larger $r^{\prime}$ may improve the decoder’s ability without benefiting the learned representations of $f$. This issue doesn’t arise in DinoV2, as the Dino-iBOT shared head is fully trained since it isn’t provided by Oquab et al. (2024).
@@ B.5 Impact of ViT backbone size
We also test the impact of the ViT backbone for ExPLoRA, varying the architecture for DinoV2 from ViT-B (86M, $\tt{L}=12$ layers, embedding dimension 768), ViT-L (303M parameters, $\tt{L}=24$ layers, embedding dimension 1024), and ViT-G (1100M parameters, $\tt{L}=40$ layers, embedding dimension 1280) for extended pre-training on fMoW-RGB. The ExPLoRA models we compare against are D-[12]-$r64$ for ViT-B, D-[24]-$r64$ for ViT-L, and D-[32]-$r32$ for ViT-G. We unfreeze the 12th, 24th, and 32nd layers for each of ViT-B, ViT-L, and ViT-G, picking these layers by extending the analysis from section˜7 to ViT-B and ViT-G. We find that the 12th (last layer) for ViT-B and the 32nd (out of 40) layer for ViT-G output representations with low mean eigenvalues compared to other layers, thus presenting good candidates for unfreezing.
In table˜15, we see that as expected, ViT-G performs the best, but is only $\uparrow 0.31\%$ better in top 1 accuracy compared to ViT-L, while using many more parameters. On the other hand, we see the highest impact for ExPLoRA on ViT-B, where the top 1 accuracy improves by $\uparrow 9.21\%$ over the original DinoV2 ViT-B. These results further demonstrate the effectiveness and efficiency of ExPLoRA as a powerful technique to create unsupervised foundation models for new visual domains.